OpenStudy (anonymous):

At what temperature will the root-mean-square speed of oxygen molecules have the value of 640 m/s? 1 kilomole of oxygen has a mass of 32kg.

OpenStudy (anonymous):

\[Vrms ^{2}=3RT/M\]

OpenStudy (anonymous):

where R= 8314 J Kmol^-1 K^-1 M= 32 Kg/Kmol

OpenStudy (anonymous):

@Rizki_Zhou what does R mean and its value

OpenStudy (anonymous):

R is universal gas constant

OpenStudy (anonymous):

The gas constant (also known as the molar, universal, or ideal gas constant, denoted by the symbol R or R) is a physical constant which is featured in many fundamental equations in the physical sciences, such as the ideal gas law and the Nernst equation. You can read more in here http://en.wikipedia.org/wiki/Gas_constant

OpenStudy (anonymous):

@Rizki_Zhou is it 640 = 3 x 8314 x T / 32 ?

OpenStudy (anonymous):

you were supposed to took \[640^{2}\], based on the equation.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

v =sqrt(3RT/m) where m is molar mass (in out case 0.032 kg) R is gas constant = 8.3 hence we have T = v^2*m/(3R) =640^2*0.032/(3*8.3) = 526.4 K = 253.25 C